OpenStudy (anonymous):

lim 2-cos3x-cos4x/x x>0 using squeezing theorem

OpenStudy (callisto):

Is this question? \[\lim_{x \rightarrow 0}\frac{2-cos3x-cos4x}{x}\]

OpenStudy (anonymous):

hey r u there..

OpenStudy (anonymous):

@Callisto yup i agree with ur question

OpenStudy (anonymous):

use ... 2-cos3x-cos4x=(1-co3x)+(1-cos4x)= 2(sin3x/2)^2 +2 (sin2x)^2

OpenStudy (anonymous):

callisto : yes, it is .

OpenStudy (anonymous):

how it can be like that, @matricked ? 2(sin3x/2)^2+2(sin2x)^2 ??